OpenStudy (anonymous):

How to solve this? lim x approache 1 (xe^x-e^x)/x^2-1

OpenStudy (turingtest):

Factor the bottom as difference of squares. Factor the numerator. Something will cancel.

OpenStudy (anonymous):

\[ \lim_{x \rightarrow 1} (xe ^{x}-e ^{x})\div[(x ^{2})-1]\]

OpenStudy (turingtest):

though L'hospital's rule is harder to use here than my method

OpenStudy (anonymous):

yeah, it's because I wrote e^1/x instead of e^1/2, but the result is right

OpenStudy (turingtest):

I would not write the 1, just e/2

OpenStudy (anonymous):

\[\lim_{x \rightarrow 1}[xe^x+e^x−e^x]/2x = \lim_{x \rightarrow 1}xe^x/2x = \lim_{x \rightarrow 1}e^x/2 = e^1/2=e/2\]

OpenStudy (turingtest):

So there you have it, two ways to solve this problem :)

OpenStudy (anonymous):

done

OpenStudy (anonymous):

the answer must be e/2. did you get it? Thanks --- Terima kasih(in malay)

OpenStudy (turingtest):

Yes, you can get that by either factoring, or using L'hospital's rule. Leebaka has done all the work for the L'hospital method.

OpenStudy (anonymous):

Sorry.. I still do not understand that solution Leebaka.. Please help me :-(

OpenStudy (anonymous):

What do you still don't understand?

OpenStudy (turingtest):

Do you know L'hospital's rule?

OpenStudy (anonymous):

I just derived the top and bottom of the function and then substituted the x for 1

OpenStudy (turingtest):

if L'hospital's rule is confusing you try my way: Factor the numerator and denominator, look for something to cancel (it will)

OpenStudy (anonymous):

I did not know L'hospital rule. Can you give that rule?

OpenStudy (anonymous):

Just use Lim f'(x)/g'(x) instead of Lim f(x)/g(x). You just need to use the derivative of the numerator and denominator, in more complex cases it is more useful than using factoration.

OpenStudy (anonymous):

If you haven't done it using factoration yet: \[\lim_{x \rightarrow 1}(xe^x-e^x)/(x^2-1) = \lim_{x \rightarrow 1} e^x(x-1)/(x+1)(x-1) = \lim_{x \rightarrow 1}e^x/(x+1) = e^1/(1+1) = e/2\]